A swimming pool is 30 yrs by 10 yrds and is 5 ft deep. How many square feet are in the surface area of the sides and bottom of the pool? It has no top.
Answer by tvandenberg(45)

Lets start by defining a few variables,
L = 30 yards = 90 ft
W = 10 yards = 30 ft
H = 5 ft
The bottom is L*W = 90 * 30 = 2700 ft^2
The shorter sides are both W*H = 30 * 5 = 150 ft^2
And the long sides are both L*H = 90 * 5 = 450 ft^2
Hope this helps.